OpenStudy (anonymous):

Find the general solution of the differential equation y - y^(4) = (y^(3) + 3x)y'

OpenStudy (anonymous):

I got an integrating factor of 1/(y^4)

OpenStudy (anonymous):

I got the answer to be -y^4-y^4+y-3xy=0

OpenStudy (anonymous):

what was your integrating factor?

OpenStudy (anonymous):

Hold up, I'm getting too many questions.

terenzreignz (terenzreignz):

@xjakester that was a y' on the right-side

OpenStudy (anonymous):

i got an answer of C = x(y^(-3)) -x - ln(y), with ln(y) being the constant i found

OpenStudy (anonymous):

it is incorrect though

OpenStudy (anonymous):

i believe the integrating factor is correct though, since the partial deriv. with respect to y of (y^(-3) -1) and the partial deriv with respect to x of (-y^(-1) - 3xy^(-4)) are equal to one another

OpenStudy (anonymous):

Are you looking for an Alternative form, or solutions?

OpenStudy (anonymous):

a general solution to an exact differential quation

terenzreignz (terenzreignz):

It's exact? Hang on... \[\Large y - y^4 = (y^3+3x)\color{green}{y'}\]

OpenStudy (anonymous):

Solving for y or x?

terenzreignz (terenzreignz):

\[\Large y - y^4 = (y^3+3x)\frac{\color{blue}{dy}}{\color{red}{dx}}\]

OpenStudy (anonymous):

its not exact initially, it is exact after you multiply it by the integrating factor y^(-4)

terenzreignz (terenzreignz):

@xjakester not really... it's calculus ^_^

OpenStudy (anonymous):

I skipped calculus for statistic. x_x

terenzreignz (terenzreignz):

I see... allow me to prod it a bit ... see if I can get something... \[\Large (y-y^4)\color{red}{dx}-(y^3+3x)\color{blue}{dy}=0\]

OpenStudy (anonymous):